OpenStudy (sphott51):

Solve( x+1)/ (3x-1) =( x+4)/(2x+6) The answer is {-5, 2} Just need to show the work

OpenStudy (anonymous):

as the math teachers say "cross multiply" and get \[(x+1)(2x+6)=(3x-1)(x+4)\] then you still have a raft of work to do to solve that quadratic equation

OpenStudy (sphott51):

It doesn't seems I got the right answer when it's suppose yo be (-5,2)

OpenStudy (anonymous):

it is in fact \(-5\) and \(2\)

OpenStudy (sphott51):

Well how do you get there?

OpenStudy (sphott51):

I did the foil in the cross multiply and combined the like-terms

OpenStudy (anonymous):

how did i get it?

OpenStudy (sphott51):

And I did the ________ x_______=12 and ______+______=8 _______x______=-12 and ______+_______=11

OpenStudy (anonymous):

\[2 x^2+8 x+6 = 3 x^2+11 x-4\]

OpenStudy (anonymous):

\[x^2+3 x-10 = 0\]

OpenStudy (sphott51):

Yeah, I got the foil

OpenStudy (anonymous):

factor and done

OpenStudy (sphott51):

How did you get x^2+3x-10=0?

OpenStudy (anonymous):

\[2 x^2+8 x+6 = 3 x^2+11 x-4\] it is a quadratic, you need to set it equal to zero subtract all terms from the left hand side to get \[0=x^2+3x-10\]

OpenStudy (sphott51):

Ohh, Okay thanks!

OpenStudy (anonymous):

yw